OpenStudy (anonymous):

verify the identity 1 + (sec^2)x *(sin^2)x = (sec^2)x

OpenStudy (anonymous):

For these types of problems it helps to write it all in terms of sin and cos. so sec x = 1/cos x, csc x = 1/ sinx, and cot x = cos x/sin x. Try this and see what you see

OpenStudy (anonymous):

\[1 + {1 \over \cos^2x} \times \sin^2x = {1 \over \cos^2x}\]

OpenStudy (anonymous):

i still am confused...this isnt my strongest subject

OpenStudy (anonymous):

\[1+\tan^2x=\sec^2x\]QED

OpenStudy (anonymous):

Follow this http://www.sosmath.com/trig/Trig5/trig5/trig5.html